A particle has initial velocity =3i+4j and a constant force F=4i-3j acts on a particle the path of the particle is
1-straight line
2-parabolic
3-circular
4-elliptical

Grade:12

2 Answers

vrun
10 years ago
Here first we will find dot product. If a vector X = Ai+Bj and Y =Ci+Dj, dot product of X and Y = AC + BD so 3*4-4*3 = V*F*cosx So x is 90 This means velocity vector is perpendicular to accelaration and here accelaration is constant. This is the case of parabolic path when it is thrown horizontaly Hence it is parabolic.
vrun
10 years ago
I hope it helps